How to write this in excel?

SSE n- 1592112 48-2 1592112 46 = V3461 1、13043
Answer #1

Answer

We need to put data for SSE and n values in two different cells

Suppose we put SSE value in cell A1, so A1 = 1592112

and n value in cell B1, so B1 = 48

then, we will calculate the S_{yx} in cell C1

We need to use sqrt formula in excel

We need to type =sqrt function in cell C1

then select A1 and B1 cells, i.e. we will have the following equation =sqrt(A1/(B1-2))

this will give us the required result

=sqrt(A1/(B1-2)) = sqrt(1592112/(48-2)) = 186.041

This is the excel result for the above problem.